# Gromov-Witten theory of $\mathsf{Hilb}^{n}(\mathbb{C}^{2})$ and Noether-Lefschetz theory of $\mathcal{A}_{g}$

### v1

@@ 4.4. Proof of Theorem 5
By Theorem 1 and the discussion in Section 4.3, $\delta_{l}\in\mathbb{Q}(t_{1},t_{2})(q)$ can be explicitly calculated.
@@
$q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}=\sum_{l}\mathsf{R}_{il}^{1}\,\big{|}_{\mathsf{t}=0}\cdot\left(\nabla_{D}u_{l}\,\big{|}_{\mathsf{t}=0}-\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}\right)\cdot\mathsf{R}_{li}^{1}\,\big{|}_{\mathsf{t}=0}\,,$
where the off-diagonal terms $\mathsf{R}_{il}^{1}\,\big{|}_{\mathsf{t}=0}$ are computed by the equation
@@
where $\nabla_{D}U$ is the diagonal matrix with diagonal entries $\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}$ and $\Psi$ is the matrix whose columns are normalized eigenvectors. The functions $q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$ are therefore rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}|_{\mathsf{t}=0}$ with coefficients in the field $\mathbb{Q}(t_{1},t_{2})(q)$, where the matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}(q)$ lie. The higher order derivatives
@@
are rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}|_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$.
@@
The matrix $\mathsf{W}$ is the Wronskian of the eigenvalues $e_{i}\,|_{\mathsf{t}=0}$ as functions of $\log(q)$. Since the eigenvalues are analytic functions of $\log(q)$, a result of Boecher (see [49, Lemma 1.12]) implies that the eigenvalues are linearly independent over $\mathbb{Q}(t_{1},t_{2})$ if and only if $\det(\mathsf{W})$ is not identically 00.
@@
$\mathsf{W}^{-1}\mathsf{A}=(\mathsf{I}_{|\mathsf{Part}(n)|\times|\mathsf{Part}(n)|}\,|\,\mathsf{W}^{-1}\widehat{\mathsf{W}})\,.$
@@
and rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$. ∎
@@ 4.5. Proof of Theorem 6
with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$ are equivariant under permutations of the indices. Therefore, after substitution in
@@
the series $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ are symmetric rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2})(q)$. After an application of Proposition 14 of Appendix A, the series $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ can be effectively reconstructed from $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ and $\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(q)$. ∎
@@ 5. Calculations
$\langle\,\mu\,\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\in\mathbb{Q}(t_{1},t_{2})(q)\,.$
@@
$\begin{split}&\langle(1,1,1,1)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=-\frac{1}{24}\cdot\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\frac{109}{12}\,,\\[5.0pt] &\langle\,(2,1,1)\,\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{35q^{5}-28q^{4}+23q^{3}+23q^{2}-28q+35}{2(q-1)(q^{2}+1)(q^{2}-q+1)}\,,\end{split}$
$\begin{split}\langle(2,2)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{-(t_{1}+t_{2})}{t_{1}t_{2}}\cdot&\frac{3}{16(q^{2}+1)^{2}(q-1)^{2}}\cdot\\ &\ \left((t_{1}^{2}+\frac{1}{18}t_{1}t_{2}+t_{2}^{2})(q^{6}+1)+(\frac{2}{9}t_{1}^{2}+5t_{1}t_{2}+\frac{2}{9}t_{2}^{2})(q^{5}+q)\right.\\ &\left.+(\frac{11}{9}t_{1}^{2}-\frac{77}{18}t_{1}t_{2}+\frac{11}{9}t_{2}^{2})(q^{4}+q^{2})+(4t_{1}^{2}+\frac{170}{9}t_{1}t_{2}+4t_{2}^{2})q^{3}\right)\,,\end{split}$
$\begin{split}\langle(3,1)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{-(t_{1}+t_{2})}{t_{1}t_{2}}\cdot&\frac{1}{2(q^{2}-q+1)^{2}(q^{2}+1)^{2}}\cdot\\[5.0pt] &\left((t_{1}^{2}+\frac{23}{36}t_{1}t_{2}+t_{2}^{2})(q^{8}+1)-(\frac{5}{6}t_{1}^{2}-\frac{41}{36}t_{1}t_{2}+\frac{5}{6}t_{2}^{2})(q^{7}+q)+(\frac{1}{3}t_{1}^{2}-\frac{257}{36}t_{1}t_{2}+\frac{1}{3}t_{2}^{2})(q^{6}+q^{2})\right.\\ &\left.+(\frac{5}{6}t_{1}^{2}+\frac{137}{12}t_{1}t_{2}+\frac{5}{6}t_{2}^{2})(q^{5}+q^{3})-(\frac{8}{3}t_{1}^{2}+\frac{170}{6}t_{1}t_{2}+\frac{8}{3}t_{2}^{2})q^{4}\right)\,,\\[5.0pt] \langle(4)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot&\frac{(q+1)}{4(q^{2}-q+1)(q^{2}+1)^{3}(q-1)}\cdot\\[5.0pt] &\left((t_{1}^{2}-\frac{5}{6}t_{1}t_{2}+t_{2}^{2})(q^{8}+1)-(\frac{5}{3}t_{1}^{2}-\frac{25}{6}t_{1}t_{2}+\frac{5}{3}t_{2}^{2})(q^{7}+q)+(2t_{1}^{2}-13t_{1}t_{2}+2t_{2}^{2})(q^{6}+q^{2})\right.\\[5.0pt] &\left.+(3t_{1}^{2}+\frac{69}{2}t_{1}t_{2}+3t_{2}^{2})(q^{5}+q^{3})-(\frac{10}{3}t_{1}^{2}+39t_{1}t_{2}+\frac{10}{3}t_{2}^{2})q^{4}\right)\,.\end{split}$
@@
$\mathsf{Tr}_{m}^{\mu}=\mathsf{trace}\left(\,\mu\,\star:QH_{\mathsf{T}}^{*}(\mathsf{Hilb}^{m}(\mathbb{C}^{2}))\to QH_{\mathsf{T}}^{*}(\mathsf{Hilb}^{m}(\mathbb{C}^{2}))\,\right)\,.$

### v2

@@ 4.4. Proof of Theorem 5
By Theorem 1 and the discussion in Section 4.3, $\delta_{l}\in\mathbb{Q}(t_{1},t_{2},q)$ can be explicitly calculated.
@@
$q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}=\sum_{l}\mathsf{R}_{il}^{1}\,\big{|}_{\mathsf{t}=0}\cdot\left(\nabla_{D}u_{l}\,\big{|}_{\mathsf{t}=0}-\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}\right)\cdot\mathsf{R}_{li}^{1}\,\big{|}_{\mathsf{t}=0}\,.$
The off-diagonal terms $\mathsf{R}_{il}^{1}\,\big{|}_{\mathsf{t}=0}$ are computed by the equation
@@
where $\nabla_{D}U$ is the diagonal matrix with diagonal entries $\nabla_{D}u_{i}\,\big{|}_{\mathsf{t}=0}$ and $\Psi$ is the matrix whose columns are normalized eigenvectors. The functions $q\frac{d}{dq}\mathsf{R}_{ii}^{1}\,\big{|}_{\mathsf{t}=0}$ are therefore rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}|_{\mathsf{t}=0}$ with coefficients in the field $\mathbb{Q}(t_{1},t_{2},q)$, where the matrix coefficients of $\mathsf{M}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{D}(q)$ lie. The higher order derivatives
@@
are rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}|_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2},q)$.
@@
The matrix $\mathsf{W}$ is the Wronskian of the eigenvalues $e_{i}\,|_{\mathsf{t}=0}$ as functions of $\log(q)$. Since the eigenvalues are analytic functions of $\log(q)$, a result of Boecher (see [51, Lemma 1.12]) implies that the eigenvalues are linearly independent over $\mathbb{Q}(t_{1},t_{2})$ if and only if $\det(\mathsf{W})$ is not identically 0.
@@
and rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2},q)$. ∎
@@ 4.5. Proof of Theorem 6
with coefficients in $\mathbb{Q}(t_{1},t_{2},q)$ are equivariant under permutations of the indices. Therefore, after substitution in
@@
the series $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ are symmetric rational functions of the $q\frac{d}{dq}$ derivatives of the eigenvalues $e_{i}\,\big{|}_{\mathsf{t}=0}$ with coefficients in $\mathbb{Q}(t_{1},t_{2},q)$. After an application of Proposition 14 of Appendix A, the series $\langle D^{\star k}\rangle^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}_{1}$ can be effectively reconstructed from $\big{\langle}D\big{\rangle}_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}$ and $\mathsf{M}_{D}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}(q)$. ∎
@@ 5. Calculations
$\langle\,\mu\,\rangle_{1}^{\mathsf{Hilb}^{n}(\mathbb{C}^{2})}\in\mathbb{Q}(t_{1},t_{2},q)\,.$
@@
$\langle(1,1,1,1)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=$ $\,-\frac{1}{24}\cdot\frac{t_{1}+t_{2}}{t_{1}t_{2}}\cdot\frac{109}{12}\,$,
$\langle\,(2,1,1)\,\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=$ $\,\frac{1}{24}\cdot\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{35q^{5}-28q^{4}+23q^{3}+23q^{2}-28q+35}{2(q-1)(q^{2}+1)(q^{2}-q+1)}\,$,
$\hskip-20.0pt\langle(2,2)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=$ $\,\frac{-(t_{1}+t_{2})}{t_{1}t_{2}}\cdot\frac{3}{16(q^{2}+1)^{2}(q-1)^{2}}\cdot$
$\ \left((t_{1}^{2}+\frac{1}{18}t_{1}t_{2}+t_{2}^{2})(q^{6}+1)+(\frac{2}{9}t_{1}^{2}+5t_{1}t_{2}+\frac{2}{9}t_{2}^{2})(q^{5}+q)\right$.
$\left.+(\frac{11}{9}t_{1}^{2}-\frac{77}{18}t_{1}t_{2}+\frac{11}{9}t_{2}^{2})(q^{4}+q^{2})+(4t_{1}^{2}+\frac{170}{9}t_{1}t_{2}+4t_{2}^{2})q^{3}\right)\,$,
$\langle(3,1)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=$ $\,\frac{-(t_{1}+t_{2})}{t_{1}t_{2}}\cdot\frac{1}{2(q^{2}-q+1)^{2}(q^{2}+1)^{2}}\cdot$
$\left((t_{1}^{2}+\frac{23}{36}t_{1}t_{2}+t_{2}^{2})(q^{8}+1)-(\frac{5}{6}t_{1}^{2}-\frac{41}{36}t_{1}t_{2}+\frac{5}{6}t_{2}^{2})(q^{7}+q)+(\frac{1}{3}t_{1}^{2}-\frac{257}{36}t_{1}t_{2}+\frac{1}{3}t_{2}^{2})(q^{6}+q^{2})\right$.
$\left.+(\frac{5}{6}t_{1}^{2}+\frac{137}{12}t_{1}t_{2}+\frac{5}{6}t_{2}^{2})(q^{5}+q^{3})-(\frac{8}{3}t_{1}^{2}+\frac{170}{6}t_{1}t_{2}+\frac{8}{3}t_{2}^{2})q^{4}\right)\,$,
$\langle(4)\rangle_{1}^{\mathsf{Hilb}^{4}(\mathbb{C}^{2})}=$ $\,\frac{(t_{1}+t_{2})^{2}}{t_{1}t_{2}}\cdot\frac{(q+1)}{4(q^{2}-q+1)(q^{2}+1)^{3}(q-1)}\cdot$
$\left((t_{1}^{2}-\frac{5}{6}t_{1}t_{2}+t_{2}^{2})(q^{8}+1)-(\frac{5}{3}t_{1}^{2}-\frac{25}{6}t_{1}t_{2}+\frac{5}{3}t_{2}^{2})(q^{7}+q)+(2t_{1}^{2}-13t_{1}t_{2}+2t_{2}^{2})(q^{6}+q^{2})\right$.
$\left.+(3t_{1}^{2}+\frac{69}{2}t_{1}t_{2}+3t_{2}^{2})(q^{5}+q^{3})-(\frac{10}{3}t_{1}^{2}+39t_{1}t_{2}+\frac{10}{3}t_{2}^{2})q^{4}\right)\,.$
@@
$\mathsf{Tr}_{m}^{\mu}=\mathsf{trace}\left(\,|\mu\rangle\,\star:QH_{\mathsf{T}}^{*}(\mathsf{Hilb}^{m}(\mathbb{C}^{2}))\to QH_{\mathsf{T}}^{*}(\mathsf{Hilb}^{m}(\mathbb{C}^{2}))\,\right)\,.$